OpenStudy (anonymous):

Help please!? How many permutations are there for the letters in the word SUCCESSES? a. 362,880 b. 22,680 c. 7560 d. 3780 my answer: a. 362,880 I got that b/c i did this 9! b/c there were 9 letters so i got 9*8*7*6*5*4*3*2*1=362,880... am i right? cuz i got kinda confused... :/

OpenStudy (anonymous):

Please explain too! :) Thank you!!! :)

OpenStudy (anonymous):

hi @jim_thompson5910 thanks for coming to help too :)

OpenStudy (lgbasallote):

successes has 9 letters yes BUT s repeats 4 times e repeats twice c repeats twice therefore it should be \[\large \frac{9!}{4!2!2!}\]

OpenStudy (anonymous):

i am confused... :( so what is the answer and why? Shouldn't it be A? what do you think @jim_thompson5910

OpenStudy (lgbasallote):

so you get \[\large \frac{9 \times 8 \times 7 \times 6 \times 5 \times \cancel{4!}}{\cancel{4!}2!2!} = \frac{9 \times 8 \times 7 \times 6 \times 5}{2 \times 2}\]

jimthompson5910 (jim_thompson5910):

It's not A. Keep in mind that you have 2 Cs. So if you simply do 9! = 362880, you're overcounting because CC is the same as CC So you need to divide by 2! = 2 to correct for this overcounting You need to do the same for the Es as well. So divide by 2! = 2 Finally, you have 4 letters that are an S. So you need to divide by 4! = 24

jimthompson5910 (jim_thompson5910):

So divide 9! = 362880 by 2*2*24 = 96 to get 362880/96 = 3780

OpenStudy (lgbasallote):

a blue owl comes to help and they ask for the gray one lol

OpenStudy (anonymous):

So the answer is d. 3780?? for sure?? 100%????

OpenStudy (lgbasallote):

yeah it's d

jimthompson5910 (jim_thompson5910):

yes I'm 100% sure

jimthompson5910 (jim_thompson5910):

If you want, you can try this idea with a smaller word that has repeating letters

OpenStudy (anonymous):

YAy! thanks so much!! Now i understand it haha it's because the letters are repeating so you have to divide by 4!2!2! ???

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

awesome! thanks! :)

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (lgbasallote):

like for example LGBASALLOTE has 11 letters and L repeats thrice and A repeats twice so you'll get \[\frac{11!}{3!2!}\] i just wanted to say my name :p

OpenStudy (anonymous):

lol! thanks! :)